# HPE-CogVLM: New Head Pose Grounding Task Exploration on Vision Language Model

### v1

@@ A.1 Details of the Designed Prompts
Table 5 provides examples of our designed prompts for HPE and BBox prediction tasks. The BBox format adheres to the specifications set by CogVLM wang2024cogvlm . Euler angles in response are first converted to positive floats. These values are then rounded to the nearest integer, formatted as strings with a fixed length of three characters, padded with zeros where necessary. This table also includes several typical examples of invalid answers to highlight how invalid answers can lead to completely ineffective outputs when multiple grounding tasks requires accurate numerical output varied in range and quantity. In response to these issues, we define a new metric described in Section 4.4 to assess the model’s availability.
A.2 Pilot Study 1: Visualization of Cross Attention Maps Supervised by Designed Prompts
Figure 4: The visualization displays cross attention maps generated in response to our custom prompts. The left image shows the attention map associated with the prompt "What is the head yaw pitch roll inside the bounding box [[335,179,445,332]]?" (bounding box for the person on the left), and the right image corresponds to the prompt "What is the head yaw pitch roll inside the bounding box [[775,105,893,261]]?" (bounding box for the person on the right).
Figure 4 provides a visual representation of cross attention maps created in response to specific prompts, which are designed to analyze HPE within designated BBoxes. The left image highlights the model’s response to the prompt of HPE task within the bounding box [[335,179,445,332]], effectively focusing on the head of the person on the left. The right image similarly demonstrates the model’s precision in targeting the head of the person on the right within the bounding box [[775,105,893,261]]. The visualizations confirm the model’s precision in focusing attention accurately on specified areas, demonstrating its localization capabilities. It also effectively demonstrates that CogVLM not only generates bounding box outputs in response to queries but also interprets bounding boxes specified within prompts.
A.3 Pilot Study 2: catastrophic Forgetting Pattern in HPE Task
@@
Table 6: The impact of catastrophic forgetting when no data rehearsal are applied.
Iterations $\boldsymbol{Acc_{test}}$ $\boldsymbol{E_{bbox}}$ $\boldsymbol{MAE_{test}}$ $\boldsymbol{E_{angle}}$
@@
Table 6 illustrates the profound impact of catastrophic forgetting in a model trained for HPE task using Agora dataset. The Refcoco test accuracy starts at a high of 91.4% at iteration 0, indicating initial proficiency in object detection. As the number of training iteration increases and the model is increasingly exposed to the HPE task, the Refcoco test accuracy drastically decreases to 10.8% at iteration 1000. This sharp decline illustrates significant forgetting of the original BBox knowledge. $E_{bbox}$ rises significantly from 0% to 36.2% at iteration 500 and then decreases to 10.2% at iteration 1000. This trend suggests that the model initially adapts to the new task at the expense of previously learned behaviors, causing a temporary increase in errors before stabilizing. The MAE transitions from "Not capable" at iteration 100 to 42.16 at iteration 1000, indicating that the model begins to gain proficiency in the new task. The decline in $E_{angle}$ from 2.5% to 0.1% implies that the model’s HPE output format becomes more consistent over time.
@@
What is particularly noteworthy in this scenario is the nature of forgetting and learning displayed by the model—old knowledge is significantly diminished before new knowledge is solidified. This is in stark contrast to human learning, where new and old knowledge often coexist and can even support the acquisition of each other. In human cognition, learning new tasks frequently involves integrating new information with existing knowledge, without the catastrophic forgetting seen in this model.
A.4 Detailed results of the influence of rehearsal ratio on multi-task learning
Table 7: The influence of rehearsal ratio on HPE task.
Model Rehearsal Ratio Refcoco Refcoco+ Refcocog CMU Panoptic
$\boldsymbol{Acc_{test}}$ $\boldsymbol{E_{bbox}}$ $\boldsymbol{Acc_{test}}$ $\boldsymbol{E_{bbox}}$ $\boldsymbol{Acc_{test}}$ $\boldsymbol{E_{bbox}}$ $\boldsymbol{MAE_{test}}$ $\boldsymbol{E_{angle}}$
Non-merging CogVLM 10% 91.1% 0% 85.2% 0% 88.9% 0% 8.18 0.13%
Non-merging CogVLM 25% 91.2% 0% 86.4% 0% 89.1% 0% 9.87 0.43%
TA merging CogVLM 10% 89.5% 0% 82.3% 0% 86.1% 0% 7.72 68.9%
TA merging CogVLM 25% 90.8% 0% 84.6% 0% 87.7% 0% 8.10 67.7%
HPE-CogVLM 10% 90.5% 0% 84.7% 0% 87.8% 0% 7.36 0.052%
HPE-CogVLM 25% 90.9% 0% 85.6% 0% 88.1% 0% 8.36 0.26%
Table 7 presents the detailed comparative results for BBox prediction task and HPE task under different rehearsal ratios. The results of Refcoco BBox and HPE MAE results have been analyzed in section 5.4. The Refcoco+ and Refcocog follow the same trend as the Refcoco dataset result, which high rehearsal ratio leads to higher BBox Prediction accuracy. But at the same time, the error ratio $E_{angle}$ also increases. This aligns with expectations, a higher rehearsal ratio helps to improve BBox predictions, yet it may reduce the importance of the new task.

### v2

@@ IV-A HPE Task Prompt Design
In some traditional CNN-based models, such as 6DRepNet and HopeNet, cropping the human head region is required as the initial step. In this paper, a new prompt method is proposed, allowing us to train HPE task utilizing the information of full images. In our prompts, BBox coordinates are leveraged to specify the human head of interest when multiple people are present. Therefore, the system is capable of effectively focusing on specific heads, which makes it easier to reduce the need for labor-intensive manual annotations and automate the inference process. Meanwhile, the global features from self-attention and head of interest features from cross-attention are both learnt to improve the robustness of HPE task. Figure 1 (b) illustrates a sample of our custom-designed prompts and responses tailored for the HPE task, demonstrating how the system interprets and responds to specific queries. Additionally, table II provides more examples of prompts and responses for the HPE and BBox prediction tasks. The BBox format adheres to the specifications set by CogVLM [26]. Euler angles in the output are first converted to positive floats. These values are then rounded to the nearest integer, formatted as strings with a fixed length of three characters, padded with zeros where necessary. This table also includes several typical examples of invalid answers to highlight how invalid answers can lead to completely ineffective outputs when multiple grounding tasks requires accurate numerical output varied in range and quantity. In response to these issues, we define a new metric described in Section IV-D to assess the model’s availability.
IV-B Datasets
Table I outlines datasets used in various stages of our framework. The CrowdHuman dataset [57] serves as the pre-training dataset due to its extensive collection of human images. Its head pose annotations are derived from pseudo-labels inferred by the pre-trained 6DRepNet [15] model, and thus are referred as weak label images. It enables the model to obtain the capability of detecting real human heads and warm up the HPE task in stage 1. The synthetic Agora dataset [21] serves as the task-specific HPE images, which encompasses full-range of human head yaw angle images and provides the GT of SMPL-X parameters [59]. Its head pose annotations are generated using the method of DirectMHP [20] [^2] https://github.com/hnuzhy/DirectMHP. This dataset is used in stages 2 and 4 of our framework to enhance HPE accuracy. The Refcoco [28], Refcoco+ [28], and Refcocog [58] train datasets, which are originally utilized by CogVLM to learn BBox prediction, are chosen as rehearsal images to help mitigate the catastrophic forgetting of existing BBox capability. In our experiments, various portions of the rehearsal images are applied to determine the optimal rehearsal ratio [29, 30].
For evaluation, a subset of the CMU Panoptic dataset serves as the test dataset for evaluating HPE task, as its panoptic images of real people closely mirror real-life scenarios. The selection of images and labels follows the DirectMHP approach [20]. To evaluate object BBox localization, the test datasets including testA and testB data from Refcoco and Refcoco+, as well as the test dataset from Refcocog, are selected as the BBox evaluation datasets.
IV-C Implementation Details
[…]
IV-D Evaluation Metrics
Angle Error Ratio ($\text{E}_{\text{angle}}$): $\text{E}_{\text{angle}}$ = $\frac{e_{\text{angle}}}{t_{\text{angle}}}$, where $e_{\text{angle}}$ denotes the number of invalid HPE answers and $t_{\text{angle}}$ denotes the number of total HPE answers. This new metric is defined to assess the capability of models to provide relevant numerical outputs for HPE task. When we prompt with a HPE query, the CogVLM could produce irrelevant responses such as an natural language processing (NLP) task response like “a person head”, a BBox task response like “[[111,222,333,444]]”, or a blended response like “[[111,999,999,99}” as shown in Table II.
BBox Error Ratio ($\text{E}_{\text{bbox}}$): $\text{E}_{\text{bbox}}$ = $\frac{e_{\text{bbox}}}{t_{\text{bbox}}}$, where $e_{\text{bbox}}$ denotes the number of invalid BBox answers and $t_{\text{bbox}}$ denotes the number of total BBox answers. This new metric is defined to assess the capability of models to provide relevant numerical outputs for BBox prediction task.
BBox accuracy (ACC.): Acc. = $\frac{m}{\hat{m}}$, where $m$ denotes the number of valid BBox answers with $\text{IoU}>0.5$ and $\hat{m}$ denotes the number of total valid BBox answers. A BBox prediction is considered to be accurate if the intersection over union (IoU) between the GT and the prediction exceeds 0.5 [28]. And the invalid answers are excluded from accuracy and MAE calculation.
MAE of Euler angles (MAE): For the HPE task, the MAE between the GT Euler angles and the predicted Euler angles is defined as follows:
where $\hat{A}_{i}$ represents the GT’s Euler angles, $A_{i}$ represents the predicted Euler angles, and variable $n$ denotes the number of valid HPE answers. The MAE is measured in a circular manner rather than linearly, leading to the inclusion of a term that minimizes the difference between the predicted and actual angle by considering a full 360-degree rotation [15, 20]. In this paper, the MAE value is considered as the average of MAE for yaw, pitch and roll Euler angles.
IV-E Baseline Methods
Traditional CNN-based models, including 6DRepNet, HopeNet and WHENet, serve as the baselines for CNN-based approaches. The 6DRepNet model, recognized as the current SOTA, is specifically retrained and tested on the same Agora and CMU datasets used in the VLM experiment to ensure a fair comparison. This model is trained 100 epochs, and the best MAE is selected for baseline analysis with our HPE-CogVLM. The pre-trained models of HopeNet and WHENet are utilized because HopeNet scripts are hard-coded and WHENet training scripts are not publicly available.
Non-merging CogVLM, directly LoRA fine-tuned model without applying model merging technique, serves as a comparison point to evaluate the effectiveness of our merging approach versus the LoRA fine-tuning only method [29, 30, 31]. The difference of Non-merging CogVLM and our HPE-CogVLM methods is that the Non-merging CogVLM bypasses stages 2 and 3, instead it undergoes significantly more training iterations in stage 4 which is equal to the total iterations of stages 2 and 4 in the HPE-CogVLM framework. For examples, our HPE-CogVLM is fine-tuned 25k and 5k iterations in stages 2 and 4 respectively, while the Non-merging CogVLM is solely fine-tuned in stage 4 for 30k iterations. This ensures fair comparison with respect to HPE task training iterations.
Task Arithmetic (TA) merging CogVLM, which adheres to our framework but replacing the layer-based merging with TA based merging, is to provide a baseline for comparing our merging approach with another merging method. The TA merging process is chosen as it forms the foundation for many other merging algorithms [51, 52]. In this process, we set the lambda parameter of task arithmetic to 0.5 [47], assigning equal importance to both the BBox prediction task and the HPE task.
V Experimental Results
V-A Baseline Comparison
TABLE III: Comparison of HPE-CogVLM performance with various baselines. The best results are highlighted in bold.
Model Refcoco Refcoco+ Refcocog CMU Panoptic
$\text{Acc}_{\text{test}}$ $\text{E}_{\text{bbox}}$ $\text{Acc}_{\text{test}}$ $\text{E}_{\text{bbox}}$ $\text{Acc}_{\text{test}}$ $\text{E}_{\text{bbox}}$ $\text{MAE}_{\text{test}}$ $\text{E}_{\text{angle}}$
WHENet - - - - - - 29.55 -
HopeNet - - - - - - 22.16 -
6DRepNet - - - - - - 10.74 -
Original Grounding CogVLM 91.4% 0% 86.7% 0% 90.2% 0% - -
Non-merging CogVLM 91.1% 0% 85.2% 0% 88.9% 0% 8.18 0.13%
TA merging CogVLM 89.5% 0% 82.3% 0% 86.1% 0% 7.72 68.9%
HPE-CogVLM 90.5% 0% 84.7% 0% 87.8% 0% 7.36 0.052%
Note: A dash (“-”) indicates that the model is either not capable of performing the specified task or not applicable to the specified metric. For instance, WHENet, HopeNet, and 6DRepNet only accept detected human head bounding boxes for HPE prediction, meaning they cannot perform bounding box detection. As a result, their performance on the Refcoco, Refcoco+, and Refcocog tasks is marked with a “-”. This applies to all the tables in the following results subsections.
The results in Table III show the performance comparison between our HPE-CogVLM and various baselines described in Section IV-E. In comparison with traditional CNN-based models, our HPE-CogVLM presents the significantly lower MAE. The HPE-CogVLM MAE is 75.1%, 66.8%, and 31.5% lower than WHENet, HopeNet, and 6DRepNet, respectively. The CNN-based models also perform worse than other CogVLM-based models, which highlights the superior robustness of VLM-based models over CNN-based models.
In comparison with Non-merging CogVLM, HPE-CogVLM MAE is 10% lower than the Non-merging CogVLM. Meanwhile the $E_{\text{angle}}$ of our model is 2.5 times smaller than the Non-merging CogVLM. This indicates that our LoRA layer-based merging method is more proficient in HPE than the method that does not utilize any model merging technique. Regarding the BBox results, the HPE-CogVLM’s BBox prediction accuracy in test datasets is 0.6%, 0.5% and 1.1% lower than the Non-merging CogVLM, however, the Non-merging CogVLM costs five times more iterations for training rehearsal images as discussed in Section IV-E. This demonstrates that even with only 1/5 of the rehearsal image training iterations, our HPE-CogVLM achieves a comparable level of BBox accuracy.
Comparing with TA merging CogVLM, our HPE-CogVLM wins in all the metrics. For instance, when evaluated on test datasets, the BBox prediction accuracy of the HPE-CogVLM exceeds that of the TA merging CogVLM by 1%, 2.4%, and 1.7%, respectively. For the HPE task performance, $E_{\text{angle}}$ of TA merging CogVLM is 68.9%, which is 1325 times larger than that of HPE-CogVLM, indicating that only 31.1% of the responses for the HPE task are valid. Due to the high number of invalid HPE responses, the MAE metric becomes ineffective for assessing the performance. This highlights that even with an additional round of fine-tuning in stage 4, the TA merging fails to produce relevant numerical responses within our research domain, ultimately proving ineffective for the HPE task. In contrast, by applying our LoRA layer-based merging method, HPE-CogVLM successfully achieves the lowest MAE and invalid output ratio, demonstrating the superiority of this approach.
V-B Catastrophic Forgetting Pattern in HPE Task
@@
TABLE IV: The impact of catastrophic forgetting when no data rehearsal are applied.
Iterations $\text{Acc}_{\text{test}}$ $\text{E}_{\text{bbox}}$ $\text{MAE}_{\text{test}}$ $\text{E}_{\text{angle}}$
@@
Table IV illustrates the profound impact of catastrophic forgetting in a model trained for HPE task only using Agora dataset. The Refcoco test accuracy starts at a high of 91.4% at iteration 0, indicating initial proficiency in object detection. As the number of training iteration increases and the model is increasingly exposed to the HPE task, the Refcoco test accuracy drastically decreases to 10.8% at iteration 1000. This sharp decline illustrates significant forgetting of the original BBox knowledge. $\text{E}_{\text{bbox}}$ rises significantly from 0% to 36.2% at iteration 500 and then decreases to 10.2% at iteration 1000. This trend suggests that the model initially adapts to the new task at the expense of previously learned behaviors, causing a temporary increase in errors before stabilizing. The MAE improves from ”Not capable” at iteration 100 to 42.16 at iteration 1000, indicating that the model begins to gain proficiency in the new task. The decline in $E_{\text{angle}}$ from 2.5% to 0.1% implies that the model’s HPE output format becomes more consistent over time.
@@
What is particularly noteworthy in this scenario is the nature of forgetting and learning displayed by the model—old knowledge is significantly diminished before new knowledge is solidified. This contrasts with human learning process, where new and old knowledge often coexist and can even enhance each other’s acquisition. In human cognition, learning new tasks frequently involves integrating new information with existing knowledge, without the catastrophic forgetting seen in this model.
V-C Selecting Optimal Rehearsal Ratios for Mitigating the Catastrophic Forgetting Problem
TABLE V: Performance of weak label CogVLM under various rehearsal ratios.
Iterations Rehearsal Ratio $\text{Acc}_{\text{test}}$ $\text{E}_{\text{bbox}}$ $\text{MAE}_{\text{test}}$ $\text{E}_{\text{angle}}$
0k 0% 91.4% 0% - -
10k 0% 21.8% 0.026% 17.20 0.48%
10k 1% 77.5% 0.19% 21.51 0.85%
10k 10% 91.0% 0% 19.32 0.32%
10k 25% 91.5% 0% 19.92 0.23%
Table V presents the performance of weak label CogVLM across different proportions (0%, 1%, 10%, 25%) [29, 30, 31] of the rehearsal dataset in stage 1. The primary aim is to determine the appropriate data rehearsal ratio to retain old knowledge for the fine-tuning in stage 4. The Refcoco test accuracy at iteration 0 is 91.4%, indicating proficiency with the BBox prediction tasks. After the training is finished, the results demonstrate a clear trend that as the rehearsal ratio increases, the Refcoco test accuracy substantially improves. Starting at a low of 21.8% when no Refcoco data is used, the accuracy spikes to 77.5% with just 1% of rehearsal ratio, eventually reaching over 91% with 10% and 25% of rehearsal ratio. This clearly shows that the more original task data used in learning a new task, the less catastrophic forgetting occurs. In the $\text{E}_{\text{bbox}}$ column, the consistently low $\text{E}_{\text{bbox}}$ values suggest that the availability of BBox predictions tend to stabilize after 10K iterations. MAE and $E_{\text{angle}}$ for HPE task show a fluctuating trend. Since the head pose weak label is provided for this pre-training stage, they may not fully reflect the model’s true HPE performance. Rehearsal ratios of 10% and 25% are selected for the stage 4 experiment due to high refcoco BBox prediction accuracy. These ratios are significantly higher than the commonly used 1% rehearsal ratio in non-grounding tasks [29, 30].
V-D The Influence of Rehearsal Ratios on Multiple grounding task Learning
(a) The influence of rehearsal ratio on MAE.
(b) The influence of rehearsal ratio on BBox Prediction.
Figure 3: The model performance under various rehearsal ratios (10% and 25%). (a) shows the MAE results under rehearsal ratio 10% and 25% on VLMs. (b) shows the Refcoco Test BBox accuracy results under rehearsal ratio 10% and 25% on VLMs.
Figure 3 presents comparative results for both the BBox prediction task and the HPE task under two different rehearsal ratios in stage 4. Between the two HPE-CogVLM models, the one with a lower rehearsal ratio (10%) achieves a better MAE of 7.36, which is 12% lower than the 8.36 observed with the higher (25%) ratio. Conversely, the Refcoco test accuracy improves slightly with higher rehearsal ratios, showing increases of 0.3% compared to the lower ratio. The similar phenomenon also presents in Non-merging CogVLM and TA merging CogVLM results. The results clearly show that a higher rehearsal ratio helps retain existing knowledge by incorporating more data from previous tasks into fine-tuning, but this comes at the cost of new task performance. So we seek for balance between the retention of old knowledge and the performance on new tasks. In our case, the 10% rehearsal ratio achieves significantly better HPE performance, while the BBox prediction is only slightly better with the 25% rehearsal ratio. After balancing both factors, the HPE-CogVLM trained with the 10% rehearsal ratio is chosen as the optimal model.
V-E Performance of HPE-oriented CogVLM on HPE Task Only
TABLE VI: The HPE-oriented CogVLM model exhibits the highest HPE performance within our framework. The best results are highlighted in bold.
Model Epochs $\text{Acc}_{\text{test}}$ $\text{MAE}_{\text{test}}$ $\text{MAE}_{\text{train}}$ $\text{E}_{\text{angle}}$
6DRepNet 3 - 12.70 9.40 -
6DRepNet 6 - 12.76 8.80 -
6DRepNet 9 - 11.44 7.90 -
6DRepNet 50 - 11.37 2.91 -
6DRepNet 100 - 11.4 2.23 -
HPE-oriented CogVLM 3 8.8% 6.40 - 0.0092%
HPE-oriented CogVLM 6 12.6% 6.31 - 0%
HPE-oriented CogVLM 9 11.0% 6.24 - 0%
In our framework, The HPE-oriented CogVLM from stage 2 stands out as the most effective model dedicated solely to the HPE task. Table VI presents comparative performance results of 6DRepNet and the HPE-oriented CogVLM, both not accommodate BBox prediction capabilities, over similar training epochs. The low Refcoco test accuracy of HPE-oriented CogVLM is expected, given that no data rehearsal is implemented in this stage so that the VLM can focus on the HPE task learning. In terms of MAE metric, the HPE-oriented CogVLM displays a gradual decrease in MAE from 6.4 at 3 epochs to 6.24 at 9 epochs. When compared our model with 6DRepNet, in the same epoch, our MAE shows much lower numbers than 6DRepNet. For example, in epoch 9, MAE of HPE-oriented CogVLM is 6.24 which is 45.5% lower than 6DRepNet. After extending the 6DRepNet training to 100 epochs, while its training MAE decreases from 9.40 to 2.23, the MAE on the CMU dataset does not improve, remaining stable around 11.4. This indicates that the model is over-fitting to the training dataset, with no enhancement in cross-dataset inference performance. This difference emphasizes the superior performance of VLM than traditional CNN-based models.
V-F Visualization of Cross Attention Maps Supervised by Designed Prompts
Figure 4: The visualization displays cross attention maps generated in response to our custom prompts. The left image shows the attention map associated with the prompt ”What is the head yaw pitch roll inside the bounding box [[335,179,445,332]]?” (BBox for the person on the left), and the right image corresponds to the prompt ”What is the head yaw pitch roll inside the bounding box [[775,105,893,261]]?” (BBox for the person on the right).
Figure 4 provides a visual representation of cross attention maps created in response to specific prompts when multiple people are present in the image. The left image highlights the model’s response to the prompt of HPE task within the BBox [[335,179,445,332]]. It effectively focuses on the head of the person on the left. The right image similarly demonstrates the model’s precision in targeting the head of the person on the right within the BBox [[775,105,893,261]]. These visualizations confirm the model’s accuracy in localizing attention within specified areas, demonstrating strong capabilities in spatial awareness. Moreover, they illustrate that CogVLM is capable not only of generating BBox outputs in response to queries but also of interpreting and responding to BBoxes specified within the prompts.
VI Conclusions
In this paper, we present a framework that successfully enhances the HPE task by leveraging the visual grounding capabilities of CogVLM. Through a novel merging approach that utilizes a high cosine similarity threshold and a “winner-takes-all” layer selection strategy, we effectively integrate HPE capabilities into the model while preserving the original BBox knowledge. This method not only improves prediction accuracy but also addresses the challenges of blended invalid response formats. Additionally, we mitigate catastrophic forgetting by optimizing the rehearsal ratio. Our experimental results demonstrate that HPE-CogVLM achieves a substantial 31.5% reduction in MAE compared to the current CNN-based state-of-the-art in cross-dataset evaluations. Furthermore, HPE-CogVLM consistently outperforms directly LoRA fine-tuned model and task arithmetic-based merging model in all HPE metrics, establishing it as a robust and effective solution for complex multimodal grounding tasks.